# Collapse of Dense Retrievers: Short, Early, and Literal Biases Outranking Factual Evidence

### v1

@@ Abstract
Dense retrieval models are commonly used in Information Retrieval (IR) applications, such as Retrieval-Augmented Generation (RAG). Since they often serve as the first step in these systems, their robustness is critical to avoid failures. In this work, by repurposing a relation extraction dataset (e.g. Re-DocRED), we design controlled experiments to quantify the impact of heuristic biases, such as favoring shorter documents, in retrievers like Dragon+ and Contriever. Our findings reveal significant vulnerabilities: retrievers often rely on superficial patterns like over-prioritizing document beginnings, shorter documents, repeated entities, and literal matches. Additionally, they tend to overlook whether the document contains the query’s answer. Notably, when multiple biases combine, models exhibit catastrophic performance degradation, selecting the answer-containing document in less than 3% of cases over a synthetic biased document without the answer. Furthermore, we show that these biases have direct consequences for downstream applications like RAG, where retrieval-preferred documents can mislead LLMs, resulting in a 34% performance drop than not providing any documents at all. $\boldsymbol{\ast}$ Code and benchmark dataset are available at https://huggingface.co/datasets/mohsenfayyaz/ColDeR.
@@ 1 Introduction
Figure 1: Paired t-test statistics comparing retriever scores between document pairs ($D_{1}$ vs. $D_{2}$) across five evaluation aspects. Document pairs are designed for controlled experiments shown in Table 1. Positive values indicate a retriever’s preference for the more biased document in each bias scenario, while for answer importance, they reflect a preference for answer-containing documents. The results show that retrieval biases often outweigh the impact of answer presence.
@@
Existing evaluations of retrieval models often focus on downstream task performance, as seen in benchmarks like BEIR Thakur et al. (2021), without probing the underlying behavior of retrievers. Some studies have analyzed specific issues in information retrieval (IR) models, such as position bias Coelho et al. (2024) or lexical overlap (Ram et al., 2023).
@@
Its population in 2008 was around 2,700 . Situated in a small plain at the foot of Mt. Terminillo …..
@@
Its population in 2008 was around 2,700 . Situated in a small plain at the foot of Mt. Terminillo …..
@@
Foil vs. Evide. Query: Who is the publisher of Assassin ’s Creed Unity?
@@
Table 1: Examples from our framework highlighting Evidence, Head Entity, and Tail Entity. In all cases, retrieval models favor Document 1 over Document 2, assigning higher retrieval scores accordingly. (Explained in §3.3)
In this work, we systematically study multiple biases’ impact on retrievers—both individually and in combination—for the first time. To enable fine-grained control over document structure and factual positioning, we repurpose a document-level relation extraction dataset (Re-DocRED Tan et al. (2022)).
@@
We further study combining multiple biases and reveal concerning patterns in current retriever architectures. When exposed to multiple interacting biases, even top-performing models exhibit dramatic degradation, selecting the answer-containing document over the foil document—filled with biases—less than 3% of the time. Moreover, we demonstrate that these biases can be exploited to manipulate Retrieval-Augmented Generation (RAG), causing retrievers to favor misleading or adversarially constructed documents, which misguide LLMs into using incorrect information and ultimately degrade its performance.
@@ 3.1 A Framework for Identifying Biases in Retrievers
To construct a retrieval dataset from Re-DocRED, we map each relation to a query template. For example, for the relation "educated at," we use the template "Where was {Head Entity} educated?" This transformation allows us to systematically examine how retrievers handle different types of factual queries.
The answers to these queries are the tail entities found in the evidence sentences provided by the dataset. For our analysis, we ensure that each query has a single evidence sentence ($S_{ev}$) within the original document ($S\in D_{orig}$) that contains both the head and tail entities. This constraint makes the sentence self-contained, allowing for precise control over the document structure in subsequent sections. We also introduce the notation $S^{+h}_{-t}$ for sentences in $D_{orig}$ that contain the head entity but not the tail entity, and $S^{-h}_{-t}$ for sentences that do not contain either entity. In each of the following sections, we will use this notation to construct a pair of document sets, $D_{1}$ and $D_{2}$, enabling a systematic investigation of retrieval score variations and potential biases. As a result, for each of our six analysis settings, we compile 250 queries, each with a single corresponding gold document, based on the test and validation sets of Re-DocRED.
3.2 Models Performance & Bias Discovery
@@
First, we evaluate several dense retrievers on the NQ dataset (Kwiatkowski et al., 2019), comparing their performance using nDCG@10 and Recall@10 metrics. Table 2 shows that Dragon models lead in performance, and the significant improvement of fine-tuned Contriever over its unsupervised counterpart highlights the importance of supervision and task-specific adaptation. Models also differ in pooling mechanisms, with Contriever using average pooling and others using CLS pooling. For details, refer to the appendix A.1.
@@ 3.3.1 Answer Importance
$t=\frac{\bar{d}}{SE(\bar{d})}=\frac{\text{Average Difference}}{\text{Standard Error}}$ (2)
@@ 3.3.2 Position Bias
Our analysis reveals a strong positional bias in dense retrievers, with models consistently prioritizing information at the beginning of documents. As shown in Figure 4, we conducted paired t-tests comparing retrieval scores when the evidence sentence is placed at different positions to scores when it is placed at the document’s beginning ($R(D_{i})-R(D_{1})$).
@@ 3.3.3 Literal Bias
Retrieval models should ideally recognize semantic equivalence across different surface forms of the same entity. For instance, a robust model should understand that "Gomes" and "Gomez" refer to the same person, or that "US" and "United States" represent the same entity. However, our analysis reveals that current models exhibit a strong bias toward exact literal matches rather than semantic matching.
@@ 3.3.4 Brevity Bias
Brevity bias refers to the tendency of retrievers to favor concise text, such as a single evidence sentence, over longer documents that include the same evidence alongside additional context. This bias is problematic because retrievers may favor a shorter, non-relevant document over a relevant one. We will discuss this potential hazard further in Section 3.5.
@@ 3.4 Interplay Between Bias Types
Table 4: The accuracy and paired t-test comparing a foil document (exploiting biases but lacking the answer) to a second document with evidence embedded in unrelated sentences. All retrieval models perform extremely poorly (<3% accuracy), highlighting their inability to distinguish biased distractors from genuine evidence.
@@
Table 4 presents the accuracy (proportion of times the model prefers $D_{2}$ over $D_{1}$), paired t-test statistics, and p-values. The results are striking: all models exhibit extremely poor performance, with accuracy dropping below 3%. The paired t-test statistics are highly negative across all models, indicating a consistent preference for foil documents over the correct evidence-containing ones. This outcome highlights the severity of bias interplay and its detrimental impact on model reliability. Furthermore, a sufficient number of biased documents can potentially cause the model to select all top-k documents from only biased results.
@@ 3.5 Impact on RAG
Model Poison Doc Foil Doc No Doc Evidence Doc
@@ Appendix A Appendix
A.1 Models Downstream Performance
We evaluate several dense retrievers on the Natural Questions (NQ) dataset (Kwiatkowski et al., 2019), comparing their performance using standard retrieval metrics: nDCG@10 and Recall@10111111Using BEIR framework (Thakur et al., 2021). The models differ in training objectives, datasets, and pooling mechanisms, offering a comprehensive view of their retrieval capabilities in our experimental setup. Table 2 (and A.2) summarizes the results.
@@ A.1 Models Downstream Performance
A.2 Position Bias: First vs. Last
@@
Model Accuracy Paired t-Test Statistic p-value
Dragon+ 0.0% -55.16 < 0.01
Dragon RoBERTa 0.0% -49.17 < 0.01
Contriever MSMARCO 0.0% -46.96 < 0.01
COCO-DR Base MSMARCO 0.0% -40.19 < 0.01
RetroMAE MSMARCO FT 0.0% -48.10 < 0.01
Contriever 1.2% -33.60 < 0.01

### v2

@@ Abstract
Dense retrieval models are commonly used in Information Retrieval (IR) applications, such as Retrieval-Augmented Generation (RAG). Since they often serve as the first step in these systems, their robustness is critical to avoid downstream failures. In this work, we repurpose a relation extraction dataset (e.g., Re-DocRED) to design controlled experiments that quantify the impact of heuristic biases, such as a preference for shorter documents, on retrievers like Dragon+ and Contriever. We uncover major vulnerabilities, showing retrievers favor shorter documents, early positions, repeated entities, and literal matches, all while ignoring the answer’s presence! Notably, when multiple biases combine, models exhibit catastrophic performance degradation, selecting the answer-containing document in less than 10% of cases over a synthetic biased document without the answer. Furthermore, we show that these biases have direct consequences for downstream applications like RAG, where retrieval-preferred documents can mislead LLMs, resulting in a 34% performance drop than providing no documents at all. (Dataset & $\boldsymbol{\ast}$ Code and benchmark dataset are available at https://huggingface.co/datasets/mohsenfayyaz/ColDeR.
@@ 1 Introduction
[…]
@@
Existing evaluations often focus on downstream task performance, as seen in benchmarks like BEIR Thakur et al. (2021), without probing the underlying behaviors of retrievers. Some studies have analyzed specific issues in information retrieval (IR) models, such as position bias Coelho et al. (2024) or lexical overlap (Ram et al., 2023).
@@
Its population in 2008 was around 2,700 . Situated in a small plain at the foot of …..
@@
Its population in 2008 was around 2,700 . Situated in a small plain at the foot of …..
@@
Foil vs. Evidence Query: Who is the publisher of Assassin ’s Creed Unity?
@@
Table 1: Construction of query–document pairs. From each Re-DocRED article, we select an Evidence Sentence ($S_{ev}$) that contains both the Head Entity ($h$, the relation’s subject) and Tail Entity ($t$, its object). We then create two variants: $D_{1}$, engineered to amplify one bias, and $D_{2}$, its control. (For position bias, e.g., $D_{1}$ is ranked higher than $D_{2}$ because $s_{ev}$ is placed at the top of $D_{1}$ but at the end of $D_{2}$.) Paired score differences $M(Q,D_{1})-M(Q,D_{2})$ over 250 queries provide the paired t-statistics plotted in Fig 3. Only in the Answer-Impact $D_{2}$ and Foil-vs.-Evidence $D_{1}$, $t$ is removed, making them non-relevant; in all other tests, both documents contain the answer and remain relevant. In all cases, retrievers favor $D_{1}$ over $D_{2}$.
In this work, we study multiple biases’ impact on retrievers—both individually and in combination—for the first time. To enable fine-grained control over document structure and factual positioning, we repurpose a document-level relation extraction dataset (Re-DocRED Tan et al. (2022)).
@@
We further study combining multiple biases and reveal concerning patterns in current retriever architectures. When exposed to multiple interacting biases, even top-performing models exhibit dramatic degradation, selecting the answer-containing document over the foil document—filled with biases—less than 10% of the time. Moreover, we demonstrate that these biases can be exploited to manipulate Retrieval-Augmented Generation (RAG), causing retrievers to favor misleading or adversarially constructed documents, which misguide LLMs into using incorrect information and ultimately degrade its performance.
@@ 3.1 A Framework for Identifying Biases in Retrievers
To construct a retrieval dataset from Re-DocRED, we map each relation to a query template (Templates are in Table A.5). For example, for the relation "educated at," we use the template "Where was {Head Entity} educated?" This transformation allows us to systematically examine how retrievers handle different types of factual queries.
@@
The answers to these queries are the tail entities found in the evidence sentences provided by the dataset. For our analysis, we ensure that each query has a single evidence sentence ($S_{ev}$) within the original document ($S_{ev}\in D_{orig}$) that contains both the head and tail entities. This constraint makes the sentence self-contained, allowing for precise control over the document structure in subsequent sections. We also introduce the notation $S^{+h}_{-t}$ for sentences in $D_{orig}$ that contain the head entity but not the tail entity, and $S^{-h}_{-t}$ for sentences that do not contain either entity. In each of the following sections, we will use this notation to construct a pair of document sets, $D_{1}$ and $D_{2}$, enabling a systematic investigation of retrieval score variations and potential biases. As a result, for each of our six analysis settings, we compile 250 queries, each with a single corresponding gold document, based on the test and validation sets of Re-DocRED.
3.2 Models Performance & Bias Discovery
First, we evaluate several dense retrievers on the NQ dataset (Kwiatkowski et al., 2019), comparing their performance using nDCG@10 and Recall@10 metrics. Table 2 shows that Dragon models lead in performance, and the significant improvement of fine-tuned Contriever (Contriever MSMARCO) over its unsupervised counterpart highlights the importance of supervision and task-specific adaptation. Models also differ in pooling mechanisms, with Contriever using average pooling and others using CLS pooling. For details, refer to the appendix A.2.
@@ 3.3.1 Answer Importance
$t=\frac{\bar{d}}{SE(d)}=\frac{\text{Average Difference}}{\text{Standard Error}}$ (2)
@@ 3.3.2 Position Bias
Our analysis reveals a strong positional bias in dense retrievers, with models consistently prioritizing information at the beginning of documents. As shown in Figure 4, we conducted paired t-tests comparing retrieval scores when the evidence sentence is placed at different positions to scores when it is placed at the document’s beginning ($M(Q,D_{i})-M(Q,D_{1})$).
@@ 3.3.3 Literal Bias
Retrievers should ideally recognize semantic equivalence across different surface forms of the same entity. For instance, a robust model should understand that "Gomes" and "Gomez" refer to the same person, or that "US" and "United States" represent the same entity. However, our analysis reveals that current models exhibit a strong bias toward exact literal matches rather than semantic matching.
@@ 3.3.4 Brevity Bias
Brevity bias refers to the tendency to favor concise text, such as a single evidence sentence, over longer documents that include the same evidence alongside additional context. This bias is problematic because retrievers may favor a shorter, non-relevant document over a relevant one. We will discuss this potential hazard further in §3.5.
@@ 3.4 Interplay Between Bias Types
ColBERT (v2) 7.6% -20.96 < 0.01
ReasonIR-8B 8.0% -36.92 < 0.01
Table 4: The accuracy and paired t-test comparing a foil document $D_{1}$ (exploiting biases but lacking the answer) to a second document $D_{2}$ with evidence embedded in unrelated sentences. Accuracy is the proportion of 250 example pairs where $M(Q,D_{2})>M(Q,D_{1})$. All retrieval models perform extremely poorly (<10% accuracy), highlighting their inability to distinguish biased distractors from genuine evidence. (Extra models explained in §A.1)
@@
Table 4 presents the accuracy (proportion of times the model prefers $D_{2}$ over $D_{1}$), paired t-test statistics, and p-values. The results are striking: all models exhibit extremely poor performance, with accuracy dropping below 10%. The paired t-test statistics are highly negative across all models, indicating a consistent preference for foil documents over the correct evidence-containing ones. This outcome highlights the severity of bias interplay and its detrimental impact on model reliability. Furthermore, a sufficient number of biased documents can potentially cause the model to select all top-k documents from only biased results.
@@ 3.5 Impact on RAG
Poison Foil No Evidence
Model Doc Doc Doc Doc
@@ 4 Conclusions
Acknowledgements
This research is partly supported by a National Science Foundation CAREER award #2339766 and an Amazon AGI Research Award. We thank Jia-Chen Gu for valuable discussions and feedback. We also appreciate the insights from our peers, and we are grateful to the anonymous reviewers for their constructive comments.
@@ Appendix A Appendix
A.1 More Models
Our study focuses on dense retrievers that generate a single embedding per document, ensuring efficiency. In contrast, newer models like ColBERT Khattab and Zaharia (2020) prioritize accuracy at the cost of efficiency. Specifically, ColBERT employs late interaction, requiring per-token embeddings for both queries and documents, which increases computational and storage demands, especially problematic for long documents. Despite these trade-offs, we evaluated ColBERT and ReasonIR-8B Shao et al. (2025) on the Foil dataset, and their performance results are in Table 4. Despite their higher computational cost, these models still fail on the Foil dataset, with under 9% correct document preferences.
A.2 Models Downstream Performance
We evaluate several dense retrievers on the Natural Questions (NQ) dataset (Kwiatkowski et al., 2019), comparing their performance using standard retrieval metrics: nDCG@10 and Recall@10131313Using BEIR framework (Thakur et al., 2021). The models differ in training objectives, datasets, and pooling mechanisms, offering a comprehensive view of their retrieval capabilities in our experimental setup. Table 2 (and A.3) summarizes the results.
@@ A.2 Models Downstream Performance
Model Accuracy Paired t-Test Statistic p-value
Contriever 0.4% -34.58 < 0.01
RetroMAE MSMARCO FT 0.4% -41.49 < 0.01
Contriever MSMARCO 0.8% -42.25 < 0.01
Dragon RoBERTa 0.8% -36.53 < 0.01
Dragon+ 1.2% -40.94 < 0.01
COCO-DR Base MSMARCO 2.4% -32.92 < 0.01
ColBERT (v2) 7.6% -20.96 < 0.01
ReasonIR-8B 8.0% -36.92 < 0.01
Table A.1: The accuracy and paired t-test comparing a foil document (exploiting biases but lacking the answer) to a second document with evidence embedded in unrelated sentences. All retrieval models perform extremely poorly, highlighting their inability to distinguish biased distractors from genuine evidence.
@@
colbert-ir/colbertv2.0 Khattab and Zaharia (2020)
reasonir/ReasonIR-8B Shao et al. (2025)
@@
A.3 Position Bias: First vs. Last
Relation ID Relation Name Query Template
P131 located in the administrative territorial entity Which administrative territorial entity is <head_entity> located in?
P577 publication date When was <head_entity> published?
P17 country Which country is <head_entity> associated with?
P264 record label Which record label is <head_entity> associated with?
P571 inception When was <head_entity> founded?
P361 part of What is <head_entity> a part of?
P800 notable work What is a notable work of <head_entity>?
P569 date of birth When was <head_entity> born?
P159 headquarters location Where is the headquarters of <head_entity> located?
P527 has part What are the components of <head_entity>?
P123 publisher Who is the publisher of <head_entity>?
P175 performer Who performed <head_entity>?
P449 original network What is the original network of <head_entity>?
P706 located on terrain feature Where is <head_entity> located on a terrain feature?
P580 start time When did <head_entity> start?
P740 location of formation Where was <head_entity> formed?
P27 country of citizenship Which country is <head_entity> a citizen of?
P403 mouth of the watercourse What is the mouth of the watercourse of <head_entity>?
P570 date of death When did <head_entity> die?
P136 genre What genre does <head_entity> belong to?
P576 dissolved, abolished or demolished When was <head_entity> dissolved or demolished?
P495 country of origin What is the country of origin of <head_entity>?
P19 place of birth Where was <head_entity> born?
P155 follows What precedes <head_entity>?
P400 platform What platform is <head_entity> available on?
P1344 participant of What was <head_entity> a participant of?
P3373 sibling Who is the sibling of <head_entity>?
P676 lyrics by Who wrote the lyrics for <head_entity>?
P26 spouse Who is the spouse of <head_entity>?
P58 screenwriter Who wrote the screenplay for <head_entity>?
P35 head of state Who is the head of state of <head_entity>?
P6 head of government Who is the head of government of <head_entity>?
P178 developer Who developed <head_entity>?
P279 subclass of What is <head_entity> a subclass of?
P127 owned by Who owns <head_entity>?
P156 followed by What follows <head_entity>?
P140 religion What is the religion of <head_entity>?
P607 conflict What conflict was <head_entity> part of?
P364 original language of work What is the original language of <head_entity>?
P463 member of Which organization is <head_entity> a member of?
P179 series What series is <head_entity> part of?
P176 manufacturer Who manufactured <head_entity>?
P190 sister city What is the sister city of <head_entity>?
P20 place of death Where did <head_entity> die?
P112 founded by Who founded <head_entity>?
P31 instance of What is <head_entity> an instance of?
P276 location Where is <head_entity> located?
P86 composer Who composed the music for <head_entity>?
P57 director Who directed <head_entity>?
P272 production company Which production company produced <head_entity>?
P50 author Who is the author of <head_entity>?
Table A.5: The query templates for each relation.
@@
Model Accuracy Paired t-Test Statistic p-value
Dragon+ 0.0% -55.16 < 0.01
Dragon RoBERTa 0.0% -49.17 < 0.01
Contriever MSMARCO 0.0% -46.96 < 0.01
COCO-DR Base MSMARCO 0.0% -40.19 < 0.01
RetroMAE MSMARCO FT 0.0% -48.10 < 0.01
Contriever 1.2% -33.60 < 0.01